# Multibanded Reduced Order Quadrature Techniques for Gravitational Wave Inference

### v1

@@ 4 Conclusions
We have described a change to the algorithm implemented in PyROQ, which significantly speeds up the construction of reduced bases for gravitational wave waveforms, by utilizing multibanded waveforms during the basis search.
We used our multiband ROQ construction method to construct a subsolar mass basis for the IMRPhenomXAS_NRTidalv3 waveform model. Comparing this to a ROQ basis constructed using the standard PyROQ method over identical parameter ranges, we found a $\sim 30\%$ decrease in basis size, and a $\sim 13$ and $\sim 9$ times decrease in construction time for the linear and quadratic bases, respectively. We also demonstrated our basis in a PE run, and found that the altered method had no affect on the final PE results relative to the standar PyROQ constructed bases.
The constructed bases for IMRPhenomXAS_NRTidalv3 are the first to be built in subsolar mass ranges, and demonstrate the capabilities of this new method. This will significantly reduce the computational cost for any parameter estimation of these longer-duration signals. Future work will focus on further implementing this method, particularly in constructing ROQs for long-duration waveforms for the next-generation detectors.

### v2

@@ 5 Conclusions
We have described a modification to the algorithm implemented in PyROQ, which significantly speeds up the construction of reduced bases for gravitational wave waveforms, by utilizing multibanded waveforms during the basis search.
We used our multiband ROQ construction method to construct an SSM basis for the IMRPhenomXAS_NRTidalv3 waveform model. Comparing this to a ROQ basis constructed using the standard PyROQ method over identical parameter ranges, we found a $\sim 30\%$ decrease in basis size, and a $\sim 13$ and $\sim 9$ times decrease in construction time for the linear and quadratic bases, respectively. We also demonstrated our basis in a PE run, and found that the altered method did not affect the final PE results relative to the standard PyROQ constructed bases.
The constructed bases for IMRPhenomXAS_NRTidalv3 are the first to be built in SSM ranges, and demonstrate the capabilities of this new method. This will significantly reduce the computational cost for any parameter estimation of these longer-duration signals. Future work will focus on further implementing this method, particularly in constructing ROQs for long-duration waveforms for the next-generation detectors. Further testing will also be performed to test the methods capability for precessing and higher mode waveforms.